MATH 367, spg 98, CHAPTER 2 Set I, Jan 23

Order with repetition, Permutations, Combinations, Multinomials.

1. How many numbers of 5 digits can be made using only the digits 1, 2, 3, 4, and 5?

2. How many different 7–digits licence plates are possible if the first 3 places are to be occupied by letters and the final 4 by numbers?

3. How many numbers of 5 digits can be made using all the digits 1, 2, 3, 4, 5?

4. In how many ways can the letters of the word FLOP be arranged?

5. In how many ways can the letters of the word FLUFF be arranged?

6. In how many ways can the letters of the word ROTOR be arranged?

7. In how many ways can the letters of the word PEPPER be arranged?

8. In how many ways can A, B, C, D, and E sit at a round table?

9. A committee of 3 is to be formed from a group of 20 people. How many different committees are possible?

For next class, Wed Jan 28: read chapter 2 and do problems 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13. Yes, that’s right, no class on Mon because of personal problems.